A reacts with B to form a compound AB. When the reaction occurs, 150.0 ml of water was heated from 17.1 degrees C to 19.8 degrees C. Assuming 1 mole of AB formed, calculate the delta H fusion for AB. Leave answer in kcal/mol.
1
Expert's answer
2017-04-28T14:26:34-0400
Q = cmΔT
ΔT = 2.7˚C C (water) = 1.007 kcal/kg · C
m (water) = V · ρ = 0.15 · 1 = 0.15 kg
Q = 1.007 · 0.15 · 2.7 = 0.41 kcal
Δ H fusion for AB = 0.41/1 = 0.41 kcal/mol